Solve. 6x + 4(3x – 2) = 6x + 2
Answer by fractalier(2101)

Okay, from
6x + 4(3x – 2) = 6x + 2
6x + 12x - 8 = 6x + 2
18x - 8 = 6x + 2
12x - 8 = 2
12x = 10
x = 10/12 = 5/6